40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4834     Entry time: Fri Jun 17 23:20:05 2011
Author: Koji 
Type: Update 
Category: LSC 
Subject: Some updates of the LSC screen 

Some updates of the LSC screen

- Signal amplitude monitor for the PD signals (--> glows red for more than 1000)

- Kissel Buttons for the main matrices

- Trigger display at the output of the DOF filters

- Signal amplitude monitor for the SUS LSC output (--> glows red for more than 10000)

 

ADC Over flow monitor is showing some unknown numbers (as ADCs are handled by IOPs).
I asked Joe for the investigation (and consideration for the policies)

Attachment 1: screen.png  43 kB  Uploaded Sat Jun 18 00:20:22 2011  | Hide | Hide all
screen.png
ELOG V3.1.3-